The commute time to work in the U.S. has a bell shaped distribution with a population mean of 24.4 minutes and a population standard deviation of 6.5 minutes.

Calculate the z-score corresponding to a commute time of 30.9 minutes

Calculate the z-score corresponding to a commute time of 17.9 minutes

Calculate the z-score corresponding to a commute time of 37.4 minutes

Calculate the z-score corresponding to a commute time of 11.4 minutes

Calculate the z-score corresponding to a commute time of 43.9 minutes

Calculate the z-score corresponding to a commute time of 4.9 minutes
Answer #1

Answer)

Z score is = (x - mean)/s.d

Given mean = 24.4

S.d = 6.5

A)

Z score for 30.9

Z = (30.9 - 24.4)/6.5

Z = 1

B)

Z = (17.9 - 24.4)/6.5

Z = -1

C)

Z = (37.4 - 24.4)/6.5 = 2

D)

Z = (11.4 - 24.4)/6.5 = -2

E)

Z = (43.9 - 24.4)/6.5 = 3

F)

Z = (4.9 - 24.4)/6.5 = -3